# Aligning Vision to Language: Text-Free Multimodal Knowledge Graph Construction for Enhanced LLMs Reasoning

### v1

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [10, 60, 2, 24] have demonstrated their superiority and versatility across various Natural Language Reasoning (NLR) tasks [53, 9, 48]. To enhance LLMs into the realm of multimodal reasoning, researchers [74, 66, 69, 59] have endeavored to equip these models with multimodal capabilities, as evidenced by advancements in Multimodal Large Language Models (MLLMs) such as BLIP-2 [38], GPT-4o[31], Janus-Pro[14], among others. Despite their notable progress, these models often experience hallucinations [5, 32], primarily arising from knowledge deficiencies due to incomplete or obsolete information.
@@
Fine-tuning LLMs demands prohibitive computational costs [30]. While text-based Knowledge Graphs (KGs) have partially addressed this limitation by efficient real-time updates [57, 67, 6], they are still restricted by modal isolation, which hinders cross-modal reasoning, as detailed in Appendix A. To bridge this semantic fragmentation, Multimodal Knowledge Graphs (MMKGs) have been developed as unified representational frameworks [41, 11, 36].
@@
In this paper, we propose VaLiK, short for Vision-align-to-Language integrated Knowledge Graph, a novel framework designed to empower LLMs with advanced multimodal reasoning. Unlike traditional methods that rely on text annotations for training extraction models and the knowledge construction process [49], VaLiK adopts a text-free approach to MMKGs construction. Specifically, we first employ several pretrained Vision-Language models (VLMs), designed based on Chain-of-Experts (CoE) principles [68], to convert visual inputs into image-specific textual descriptions through cross-modal feature alignment. This procedure eliminates the need for manually annotated image captions in both the knowledge extraction and construction phases while preserving visual details typically missing in generic text descriptions. Moreover, in contrast to existing relation detection methods that require predefined label taxonomies [79, 17, 55, 76], VaLiK excels at extracting profound semantic relationships that are both KG-compatible and capture novel associations beyond training supervision. While VLMs enable cross-modal reasoning and interpretation, they introduce spurious relational noise through hallucinated inter-modal attributions, as depicted in Figure 2. We address this limitation through cross-modal similarity recalibration, strategically filtering inconsistent information while preserving valid semantic correspondences. Finally, the purified descriptions are systematically organized into MMKGs via LLM-driven symbolic structuring [26], bridging visual and textual domains with factual consistency.
@@
In summary, VaLiK is the first framework that enables end-to-end, text-free, zero-shot, and storage-efficient multimodal knowledge construction with high adaptability and scalability. Our key contributions include:
  • To the best of our knowledge, VaLik is the first end-to-end framework to build Text-Free MMKGs to improve LLMs’ multimodal reasoning capabilities, effectively eliminating the need for manually annotated textual material and enabling a completely autonomous multimodal knowledge generation process.
@@ 2.2 Knowledge-Augmented Multimodal Learning
Multimodal learning has seen significant progress in aligning and integrating information across different data modalities [7, 70]. The incorporation of structured knowledge through MMKGs further enhances these approaches, improving the reasoning capabilities and generalization across a variety of domains, such as visual question answering [62, 46, 54], recommendation systems [56, 65, 18], and classification [78, 50, 29]. Methods like GraphAdapter’s dual-KG adaptation[39] and contrastive multi-relational encoding with KGs[21] inject external knowledge into models, refining their performance and improving their capability to handle complex tasks. Additionally, Lee et al. [36] proposed MR-MKG, a novel framework that constructs task-specific MMKGs to enhance multimodal reasoning in LLMs. These knowledge-augmented paradigms demonstrate superior cross-modal semantic grounding compared to unimodal approaches [33, 15]. However, their reliance on preconstructed MMKGs often leads to domain discrepancies, where generic knowledge schemas misalign with task-specific reasoning patterns, ultimately limiting contextual precision in target applications.
@@ 2.3 Multimodal Large Language Models
The limitations of text-only LLMs in meeting increasingly complex demands have spurred extensive research [73, 80, 77] into developing LLMs capable of effectively processing and reasoning over multimodal inputs. Current research predominantly employs adapter or projection layers to connect the embedding spaces of various modality-specific encoders with the textual embedding space of LLMs [36]. For instance, foundational models like CLIP [51] and BLIP [37] pioneered cross-modal alignment by jointly training vision and text encoders to map images and text into a shared embedding space. Building on this, LLaVA [40] and Flamingo [4] advanced the field by integrating visual encoders with LLMs, enabling more nuanced multimodal understanding and generation. More recently, Gemini [58], Qwen2-VL [63] and GPT-4o [31] have further pushed the boundaries by scaling up multimodal pretraining and introducing sophisticated mechanisms for cross-modal interaction. However, multimodal LLMs remain prone to hallucinations. While they enhance cross-modal alignment, they neither acquire new knowledge nor avoid introducing noise through integration. To address these limitations, VaLiK ”uses the master’s tools to refine the master’s craft,” first constructing MMKGs via MLLMs and then leveraging them to enhance MLLMs’ reasoning capabilities.
@@ 3.3 MMKG Construction for Enhanced Reasoning
We begin by refining the generated textual description $\hat{S}$ (VLM-based information), which is then optionally concatenated with any available external textual knowledge $T$ to form the input for KG generation. This combined input is used to generate MMKGs with the help of the LightRAG model [26] for its efficiency in multi-hop reasoning and dynamic knowledge integration.
$\mathcal{G}=\text{LightRAG}\left(\hat{S}\oplus T\right),$ (9)
where $\oplus$ denotes optional concatenation based on the availability of $T$. The LightRAG architecture processes this combined input through a LLM to generate graph structures capturing both visual and textual relationships.
The resulting knowledge graph $\mathcal{G}=(\mathcal{E},\mathcal{R})$ consists of entities $\mathcal{E}$ and relations $\mathcal{R}$, which are structured as triplets. Each triplet represents a relationship between two entities. This MMKG is formally expressed as:
@@
where $\mathcal{E}$ represents the set of entities in the graph. These entities can be objects or concepts that appear in the image or external texture knowledge. $\mathcal{R}$ represents the set of relations between entities, such as ”is a type of,” ”part of,” or ”has property. A triplet $(h,r,t)$ consists of a head entity $h$, a relation $r$, and a tail entity $t$, where both $h$ and $t$ are elements from $\mathcal{E}$, and $r$ is an element from $\mathcal{R}$.
Given an input text corpus $\mathcal{T}$, LightRAG employs an extraction function $f_{\text{ERE}}$:
$f_{\text{ERE}}:\mathcal{T}\rightarrow(\mathcal{E},\mathcal{R}),\quad\mathcal{T}=\hat{S}\oplus T,$ (11)
where =$\mathcal{E}=\{e_{1},\ldots,e_{n}\}$ denotes extracted entities and =$\mathcal{R}=\{(h_{i},r_{i},t_{i})\}$ contains relation triplets.
Multimodal Reasoning Enhancement. Following LightRAG’s hybrid strategy, we retrieve relevant triplets from $\mathcal{G}$ through structural patterns during LLMs inference:
$\mathcal{G}_{q}=\text{Retrieve}_{\text{global+local}}(q,\mathcal{G}),$ (12)
where $\text{Retrieve}_{\text{global}}$ finds broad themes and topics by grouping related entities, and $\text{Retrieve}_{\text{local}}$ focuses on specific entities and their direct connections to capture detailed relationships. More explanations are provided in Appendix D.
@@ 4 Experiment
Table 1: The performance evaluation of zero-shot text-only LLMs without any fine-tuning on the training set. Since these models handle text exclusively, the test data is formatted in an unimodal text representation for compatibility. In our implementations, both LightRAG and VaLiK adopt Qwen2.5-7B as the base reasoning model. Bold indicates the highest value, and underline indicates the second highest.
@@ 4.1 Setups
Evaluation Datasets.
We evaluate VaLiK on two multimodal reasoning benchmarks with distinct characteristics:
@@
Task Formulation.
For CrisisMMD, we design three core multimodal classification tasks [^1] This setting references the repository GitHub and Abavisani et al. [1]: (1) information relevance filtering through binary classification, (2) fine-grained humanitarian category recognition, and (3) a consolidated taxonomy with merged categories to reduce label complexity. We exclude the unimodal damage assessment task from the original dataset to focus on multimodal interactions. For ScienceQA, we follow the original evaluation framework with multi-dimensional metrics: question types, contextual modalities, and educational stages. Performance is assessed via accuracy percentage across these categories.
Baselines.
We conduct a comprehensive evaluation of text-only LLMs, multimodal VLMs, and KGs that enhance LLMs in multimodal reasoning.
  • For CrisisMMD, we compare zero-shot text-only LLMs (LLaMA-2 [60], GPT-4 [2], DeepSeek-R1 [24], Qwen-2.5 [71]) and multimodal VLMs (CLIP [51], LLaVA [40], GPT-4o [31], Qwen2-VL [63], BLIP-2 [38]).
  • For ScienceQA, we compare models for general domains in zero/few-shot settings, including text-only LLMs (GPT Model [43], CoT [43], DDCoT [80]), multimodal VLMs (LG-VQA [23], LaVIN [45], BLIP-2) and Tool-LLM Chameleon [44]. These models are not specifically fine-tuned for scientific tasks, ensuring a fair evaluation of generalization capabilities.
  • We further compare the multimodal reasoning performance of LLMs assisted by KGs, evaluating text-based KGs built with LightRAG, and pre-constructed MMKGs such as Visual Genome [35] and Mmkg [41].
Implementation.
For MMKG construction, we design a chain of VLMs including BLIP-2, LLaVA, and Qwen2-VL, with the CLIP-ViT-L/14 for pruning. Stronger or additional VLMs could be employed to enhance performance if more computational resources are available. We construct MMKGs using the entire training set as our knowledge base. In comparative experiments, LightRAG utilizes only textual data, while VaLiK employs two configurations: (1) fully image-generated text descriptions (Image-only), and (2) original text combined with image-generated text (Text-Image). Dynamic window partitioning based on sentence length ensures syntactically coherent pruning results. Similarity thresholds are set to $\tau=0.25$ for CrisisMMD and $\tau=0.20$ for ScienceQA based on empirical evaluations to balance precision and recall. We construct the graph using DeepSeek-R1-70B and implement LightRAG’s hybrid retrieval approach with Qwen2.5-7B. For graph construction and multimodal reasoning, we utilize 1×NVIDIA A100-80GB GPUs. Task-specific prompts are designed to assist LLMs in multimodal reasoning evaluation.
@@ 4.2 Main Results
Table 4: Ablation study on ScienceQA benchmark (CVs: CoE-based Vision-Language Models; SV: Similarly Verification).
@@ 4.4 Further Analysis
Figure 4: VLM Quantity Impact Analysis on CrisisMMD.
@@ Appendix C Case Studies on Visual Specificity Deficits in VLM-Generated Captions
As exemplified in Figure 7, vision-language models like BLIP-2 [38] tend to produce oversimplified textual descriptions that critically lack actionable visual-semantic signals. The VLM-generated caption (”A map of the united states with the location of the united states”) merely identifies coarse-grained scene semantics, failing to capture object-level attributes (color coding of regions), spatial relationships (border adjacency between Arizona and Mexico) and compositional context (compass orientation in lower-right corner). In contrast, human annotations (”This is a map of the United States. The main part of the country is shown in green, with several states labeled. Arizona is in the southwestern part of the US, bordering Mexico. Oklahoma is in the central - southern region. Louisiana is located along the Gulf of Mexico in the southeastern part. West Virginia is in the eastern part of the country. There’s also a compass in the bottom - right corner to show directions.”) demonstrate essential characteristics for multimodal reasoning.
Appendix D LightRAG
LightRAG is an excellent project that can effectively assist us in achieving the goal of automatically constructing MMKGs[26]. However, there are still some details of LightRAG that need to be explained. Firstly, let’s talk about retrieval. LightRAG introduces text chunks during the retrieval process, so the retrieved information is more extensive. By using high-level and low-level keywords combined with text chunk vector retrieval, LightRAG can retrieve more information. Meanwhile, the retrieval model of LightRAG is also crucial. Larger LLMs have slower retrieval speeds but better performance. In this experiment, we used Qwen2.5-7B for retrieval. We also tested the retrieval performance of 32B and 72B models, which showed a 1%-5% improvement in performance, but it also significantly increased the graph construction time. Therefore, we finally adopted a lightweight retrieval model. The details of the entire LightRAG are shown in Algorithm  1.

### v2

# Aligning Vision to Language: Annotation-Free Multimodal Knowledge Graph Construction for Enhanced LLMs Reasoning

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [10, 66, 2, 26] have demonstrated their superiority and versatility across various Natural Language Reasoning (NLR) tasks [59, 9, 54, 44]. To enhance LLMs into the realm of multimodal reasoning, researchers [80, 72, 75, 65] have endeavored to equip these models with multimodal capabilities, as evidenced by advancements in Multimodal Large Language Models (MLLMs) such as BLIP-2 [41], GPT-4o [33], Janus-Pro [14], among others. Despite their notable progress, these models often experience hallucinations [5, 35], primarily arising from knowledge deficiencies due to incomplete or obsolete information.
@@
Fine-tuning LLMs demands prohibitive computational costs [32]. While text-based Knowledge Graphs (KGs) have partially addressed this limitation by efficient real-time updates [63, 73, 6], they are still restricted by modal isolation, which hinders cross-modal reasoning, as detailed in Appendix A. To bridge this semantic fragmentation, Multimodal Knowledge Graphs (MMKGs) have been developed as unified representational frameworks [34, 46, 11, 39].
@@
In this paper, we propose VaLiK, short for Vision-align-to-Language integrated Knowledge Graph, a novel framework designed to empower LLMs with advanced multimodal reasoning. Unlike traditional methods that rely on text annotations for training extraction models and the knowledge construction process [55], VaLiK adopts a annotation-free approach to MMKGs construction. Specifically, we first employ several pretrained Vision-Language models (VLMs), designed based on Chain-of-Experts (CoE) principles [74], to convert visual inputs into image-specific textual descriptions through cross-modal feature alignment. This procedure eliminates the need for manually annotated image captions in both the knowledge extraction and construction phases while preserving visual details typically missing in generic text descriptions. Moreover, in contrast to existing relation detection methods that require predefined label taxonomies [85, 17, 61, 82], VaLiK excels at extracting profound semantic relationships that are both KG-compatible and capture novel associations beyond training supervision. While VLMs enable cross-modal reasoning and interpretation, they introduce spurious relational noise through hallucinated inter-modal attributions, as depicted in Figure 2. We address this limitation through cross-modal similarity recalibration, strategically filtering inconsistent information while preserving valid semantic correspondences. Finally, the purified descriptions are systematically organized into MMKGs via LLM-driven symbolic structuring [28], bridging visual and textual domains with factual consistency.
@@
In summary, VaLiK is the first framework that enables end-to-end, annotation-free, zero-shot, and storage-efficient multimodal knowledge construction with high adaptability and scalability. Our key contributions include:
  • To the best of our knowledge, VaLik is the first end-to-end framework to build Annotation-Free MMKGs to improve LLMs’ multimodal reasoning capabilities, effectively eliminating the need for manually annotated textual material and enabling a completely autonomous multimodal knowledge generation process.
@@ 2.2 Knowledge-Augmented Multimodal Learning
Multimodal learning has seen significant progress in aligning and integrating information across different data modalities [7, 76, 45]. The incorporation of structured knowledge through MMKGs further enhances these approaches, improving the reasoning capabilities and generalization across a variety of domains, such as visual question answering [68, 51, 60], recommendation systems [62, 71, 18], and classification [84, 56, 31]. Methods like GraphAdapter’s dual-KG adaptation [42] and contrastive multi-relational encoding with KGs [23] inject external knowledge into models, refining their performance and improving their capability to handle complex tasks. Additionally, Lee et al. [39] proposed MR-MKG, a novel framework that constructs task-specific MMKGs to enhance multimodal reasoning in LLMs. These knowledge-augmented paradigms demonstrate superior cross-modal semantic grounding compared to unimodal approaches [36, 15]. However, their reliance on preconstructed MMKGs often leads to domain discrepancies, where generic knowledge schemas misalign with task-specific reasoning patterns, ultimately limiting contextual precision in target applications.
@@ 2.3 Multimodal Large Language Models
The limitations of text-only LLMs in meeting increasingly complex demands have spurred extensive research [79, 86, 83] into developing LLMs capable of effectively processing and reasoning over multimodal inputs. Current research predominantly employs adapter or projection layers to connect the embedding spaces of various modality-specific encoders with the textual embedding space of LLMs [39]. For instance, foundational models like CLIP [57] and BLIP [40] pioneered cross-modal alignment by jointly training vision and text encoders to map images and text into a shared embedding space. Building on this, LLaVA [43] and Flamingo [4] advanced the field by integrating visual encoders with LLMs, enabling more nuanced multimodal understanding and generation. More recently, Gemini [64], Qwen2-VL [69] and GPT-4o [33] have further pushed the boundaries by scaling up multimodal pretraining and introducing sophisticated mechanisms for cross-modal interaction. However, multimodal LLMs remain prone to hallucinations. While they enhance cross-modal alignment, they neither acquire new knowledge nor avoid introducing noise through integration. To address these limitations, VaLiK "uses the master’s tools to refine the master’s craft," first constructing MMKGs via MLLMs and then leveraging them to enhance MLLMs’ reasoning capabilities.
@@ 3.3 MMKG Construction for Enhanced Reasoning
We begin by refining the generated textual description $\hat{S}$ (VLM-based information), which is then optionally concatenated with any available external textual knowledge $T$ to form the input for KG generation. This combined input is used to generate MMKGs with the help of a LLM [22, 28], leveraging its capacity for multi-hop reasoning and dynamic knowledge integration.
$\mathcal{G}=\text{LLM}\left(\hat{S}\oplus T\right),$ (9)
where $\oplus$ denotes optional concatenation based on the availability of $T$. The resulting graph $\mathcal{G}$ captures both visual and textual relationships inferred by the LLM.
We define $\mathcal{G}$ as a set of triplets:
@@
where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations. Entities include objects or concepts from the image or external text, while relations describe connections such as “is a type of,” “part of,” or “has property.” Each triplet $(h,r,t)$ links a head entity $h$ and a tail entity $t$ via relation $r$.
Multimodal Reasoning Enhancement. To support multimodal reasoning, we retrieve relevant triplets from $\mathcal{G}$ through structural patterns during LLMs inference:
$\mathcal{G}_{q}=\text{Retrieve}(q,\mathcal{G}),$ (11)
where $\text{Retrieve}(\cdot)$ denotes a retrieval strategy that identifies subgraphs relevant to the query for reasoning. Detailed retrieval strategies are described in Appendix D.
@@ 4 Experiment
Table 1: The performance evaluation of text-only LLMs using few-shot prompting without any fine-tuning on the training set. As these models handle text only, test data is formatted as unimodal text for compatibility. In our implementations, both LightRAG and VaLiK adopt Qwen2.5-7B as the base reasoning model. Bold indicates the highest value, and underline indicates the second highest.
@@ 4.1 Setups
Evaluation Datasets. We evaluate VaLiK on two multimodal reasoning benchmarks with distinct characteristics:
@@
Task Formulation. For CrisisMMD, we define three multimodal classification tasks [^1] This setting references the repository GitHub and Abavisani et al. [1]: (1) binary information relevance filtering, (2) fine-grained humanitarian category recognition, and (3) a consolidated taxonomy with merged categories to reduce label complexity. We omit the unimodal damage assessment to focus on multimodal aspects. For ScienceQA, we follow the original evaluation using multiple metrics: question types, contextual modalities, and educational stages. Performance is assessed through accuracy percentage across these categories.
Baselines. We conduct a comprehensive evaluation of text-only LLMs, multimodal VLMs, and KGs that enhance LLMs in multimodal reasoning.
  • For CrisisMMD, we compare text-only LLMs using few-shot prompting (LLaMA-2 [66], GPT-4 [2], DeepSeek-R1 [26], Qwen-2.5 [77]) and multimodal VLMs (CLIP [57], LLaVA [43], GPT-4o [33], Qwen2-VL [69], BLIP-2 [41]).
  • For ScienceQA, we compare models for general domains in zero/few-shot settings, including text-only LLMs (GPT Model [48], CoT [48], DDCoT [86]), multimodal VLMs (LG-VQA [25], LaVIN [50], BLIP-2, CCOT [53], GraphVis [19]) and Tool-LLM Chameleon [49]. These models are not specifically fine-tuned for scientific tasks, ensuring a fair evaluation of generalization capabilities.
  • We further compare the multimodal reasoning performance of LLMs assisted by KGs, evaluating text-based KGs built with LightRAG [28], and pre-constructed MMKGs such as Visual Genome [38] and Mmkg [46].
Implementation. For MMKG construction, we design a chain of VLMs including BLIP-2, LLaVA, and Qwen2-VL, with the CLIP-ViT-L/14 for pruning. Stronger or additional VLMs could be employed to enhance performance if more computational resources are available. We use the entire training set as the knowledge base and construct MMKGs from the extracted descriptions based on the LightRAG framework. In comparative experiments, the LightRAG method we evaluate utilizes only textual data, while VaLiK employs two configurations: (1) fully image-generated text descriptions (Image-only), and (2) original text combined with image-generated text (Text-Image). Dynamic window partitioning based on sentence length ensures syntactically coherent pruning results. Similarity thresholds are set to $\tau=0.25$ for CrisisMMD and $\tau=0.20$ for ScienceQA based on empirical evaluations to balance precision and recall. See Appendix E for selection details. We construct the graph using DeepSeek-R1-70B and implement LightRAG’s hybrid retrieval approach with Qwen2.5-7B. For graph construction and multimodal reasoning, we utilize 1×NVIDIA A100-80GB GPUs. Task-specific prompts are designed to assist LLMs in multimodal reasoning evaluation.
@@ 4.2 Main Results
Table 4: Ablation study on ScienceQA benchmark (CVs: CoE-based Vision-Language Models; SV: Similarly Verification). Performance metrics include: NAT (natural science), SOC (social science), LAN (language), TXT (text context), IMG (image context), NO (no context), G1-6 (grades 1-6), G7-12 (grades 7-12).
@@ 4.4 Further Analysis
Computational Costs. Due to space limitations, we provide an overview of VaLiK’s computational costs in Appendix F. Our method is significantly more cost-effective than manual annotation or LLM fine-tuning.
Figure 4: Impact analysis of VLM quantity on CrisisMMD.
@@ 5 Conclusion
6 Acknowledgments
The research was supported by Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (Grant No. 2022ZD0160201) and the Science and Technology Commission of Shanghai Municipality (Grant No. 22DZ1100102).
@@ Appendix C Case Studies on Visual Specificity Deficits in VLM-Generated Captions
As exemplified in Figure 7, vision-language models like BLIP-2 [41] tend to produce oversimplified textual descriptions that critically lack actionable visual-semantic signals. The VLM-generated caption ("A map of the united states with the location of the united states") merely identifies coarse-grained scene semantics, failing to capture object-level attributes (color coding of regions), spatial relationships (border adjacency between Arizona and Mexico) and compositional context (compass orientation in lower-right corner). In contrast, human annotations ("This is a map of the United States. The main part of the country is shown in green, with several states labeled. Arizona is in the southwestern part of the US, bordering Mexico. Oklahoma is in the central - southern region. Louisiana is located along the Gulf of Mexico in the southeastern part. West Virginia is in the eastern part of the country. There’s also a compass in the bottom - right corner to show directions.") demonstrate essential characteristics for multimodal reasoning.
Appendix D Retrieval Strategy in MMKG Construction
We adopt retrieval strategies based on the framework provided by LightRAG [28], which supports multiple modes:
  • local: focuses on context-dependent information;
  • global: utilizes global knowledge;
  • hybrid: combines local and global retrieval methods;
  • naive: performs basic search without advanced techniques;
  • mix: integrates knowledge graph and vector retrieval;
In our implementation, we rely on the hybrid retrieval mode, which balances the precision of local cues with the breadth of global knowledge. This strategy improves the relevance and completeness of retrieved information, which is crucial for high-quality MMKG construction.
LightRAG is an excellent project that effectively supports automatic MMKG construction, and its retrieval design plays a central role in our framework. Specifically, LightRAG introduces keyword-guided text chunking to expand the retrievable context. By leveraging both high-level and low-level keywords in combination with chunk-level vector retrieval, it enables more comprehensive knowledge access. In addition, the choice of the retrieval model is also important. Larger LLMs have slower retrieval speeds but better performance. In this experiment, we used Qwen2.5-7B for retrieval. We also tested the retrieval performance of 32B and 72B models, which showed a 1%-5% improvement in performance, but it also significantly increased the graph construction time. Therefore, we finally adopted a lightweight retrieval model. The details of the entire LightRAG are shown in Algorithm  1.
@@ Appendix D Retrieval Strategy in MMKG Construction
Appendix E Selection of Sensitivity Threshold $\tau$
We select the sensitivity threshold $\tau$ empirically based on performance on the validation set. In practice, $\tau$ can be approximately determined by observing the token length distribution of captions: datasets with richer visual content and longer captions tend to benefit from a lower $\tau$, while simpler datasets can tolerate a higher $\tau$. This provides a practical way to adjust $\tau$ without extensive tuning.
In addition, we notice a key pattern when analyzing the relevance scores across windows. Around certain values of $\tau$, the scores tend to cluster tightly on both sides of the threshold. As a result, even a small change in $\tau$ near these points can lead to a large change in the number of tokens being pruned. This indicates that the pruning process is especially sensitive around those points, and adjusting $\tau$ even slightly may have a big impact on the final token budget.
Appendix F Construction Cost and Scalability
Construction cost is a complex issue, which we analyze from the perspectives of time and hardware requirements. Time-wise, the main components are CoE and LightRAG. While using APIs can significantly speed up the process, offline deployment and inference are also feasible. For example, generating descriptions with Qwen2-VL-7B achieves around 60 tokens per second, processing one image every 4 seconds. Thus, processing 1k images takes approximately 1.21 hours. Constructing a KG with Qwen2.5-7B yields about 196k tokens per hour, leading to a total of 1.33 hours for 1k images. The intermediate pruning step, accelerated by CLIP’s fast processing speed, is negligible. Overall, the cost is much lower than manual annotation or fine-tuning LLMs, making the method applicable to large-scale datasets. For resource-constrained users, deploying a lightweight VLM with CoE is comparable to or even more efficient than deploying a powerful VLM, further demonstrating the scalability of our approach.
Appendix G Discussion on VLM Usage and Design Flexibility
Our observations on the number and type of VLMs used in CoE are consistent with the original conclusions drawn in the CoE paper [74]. Regardless of the specific VLM architecture, increasing the number of models $N$ consistently improves performance up to a saturation point, after which further scaling yields diminishing returns. Moreover, we find that convergence is achieved more quickly when using lower softmax temperatures or simpler datasets. These factors reduce the ambiguity in model disagreement, allowing consensus to form more rapidly among the ensemble.
Interestingly, our results also show that using a single, strong VLM can achieve performance comparable to a cascade of smaller, lightweight models. This suggests a practical trade-off between model strength and ensemble size—while ensembling helps in reaching consensus across diverse weak learners, a single high-capacity model may suffice in many scenarios, especially when computational resources are limited.
In the original CoE method, the outputs from all VLM experts are first aggregated together, and then a selection process determines which expert descriptions to use. To save time in constructing the MMKGs with LLMs, we instead adopted a sequential strategy where the output of one expert is used as the prompt input for the next. We also evaluated the original aggregation and selection strategy on a smaller-scale dataset and found it to perform well, sometimes even surpassing the sequential approach. This confirms that CoE’s original design of aggregating all experts’ outputs before selecting which descriptions to use is effective and remains a strong baseline. However, correspondingly, using LLMs to construct MMKGs based on these aggregated descriptions requires significantly more time.
Additionally, while we apply pruning only at the final description step, pruning during intermediate steps may also yield good results depending on the dataset and task. There is no fixed rule for when or how to apply pruning, and our framework is designed to be flexible enough to accommodate different strategies. We emphasize that both our CoE framework and the SV step are intended to be adaptable, allowing users to experiment freely and select the approach that best suits their needs.
There are various VLMs that can be used for pruning. Among them, we recommend CLIP due to its fast inference speed and pruning performance comparable to other VLMs. Given its efficiency and effectiveness, CLIP serves as a practical choice for pruning in many scenarios.